OpenStudy (anonymous):

Solve the equation. Show your work. 8x – 12 = 4x + 24

Parth (parthkohli):

8x = 4x + 36 4x = 36 x = 36/4 x = 9

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

what the .... you give as all of your homework. try to practice by your self man.

OpenStudy (anonymous):

His just an answer smocher....

OpenStudy (cwrw238):

add 12 to both sides subtract 4x from both sides divide both sides by coefficient of x

OpenStudy (anonymous):

he will get the answer and close the question. what the... no appreciation.

Parth (parthkohli):

He's giving all the questions of his HW and that too of 6th grade!

Parth (parthkohli):

Give something tough, dude

OpenStudy (anonymous):

@ParthKohli hahaha, i learned them when i was in my mothers womb. LOL!

Parth (parthkohli):

@kumarsajan Hahaha

OpenStudy (anonymous):

@ParthKohli @kumarsajan Come on guys you are making him feel bad.